# Boosting Adversarial Transferability with Low-Cost Optimization via Maximin Expected Flatness

### v2

@@ III-C Design Challenges
Existing flatness-enhanced methods face two unresolved challenges: theoretical ambiguity and optimization imbalance. First, the heuristic adoption of flatness metrics lacks rigorous theoretical grounding to connect geometric flatness with adversarial transferability [7, 8], with some work even questioning their inherent correlation [2], leading to inconsistent metric selection and suboptimal performance. Second, as we shown in Section III-B, current methods suffer from a critical exploitation-exploration trade-off: average-case flatness optimization over-exploits local gradient consensus, causing premature convergence to suboptimal plateaus, while worst-case criteria induce over-exploration of non-informative sensitivity peaks, resulting in oscillatory updates. Addressing these challenges requires developing a unified framework that integrates theoretical insights into flatness-transferability relationships with adaptive mechanisms to harmonize local exploitation and global exploration, thereby maximizing adversarial transferability.
@@ IV Theoretical Analysis
Building upon the systematic formalization of flatness metrics and identified challenges in Section III, this section resolves the first key challenge: the theoretical ambiguity between geometric flatness and adversarial transferability. We first establish a formal definition of transferability by introducing the adversarial loss gap, then rigorously demonstrate that improved multiple-order average-case flatness enhances adversarial transferability. This theoretical foundation directly addresses the heuristic limitations of existing methods while guiding principled optimization criteria.
@@ IV-A Adversarial Transferability Formulation
To formally characterize adversarial transferability, we first define the adversarial loss gap (ALG) as the loss difference between adversarial and natural examples for a given model $F$ (Definition 2). A larger ALG indicates stronger perturbation efficacy, where the adversarial example $\mathbf{x}+\boldsymbol{\delta}$ significantly amplifies the model’s loss compared to the original input $\mathbf{x}$. This metric directly quantifies how effectively $\boldsymbol{\delta}$ destabilizes $F$’s predictions, serving as the basis for analyzing transferability.
Definition 2 (Adversarial Loss Gap).
Given a model $\mathcal{F}$, an input–label pair $\bigl{(}\mathbf{x},y\bigr{)}$, and an adversarial perturbation $\boldsymbol{\delta}$, the adversarial loss gap (ALG) is defined as:
$\operatorname{ALG}\!\bigl{(}(\mathbf{x},y),\boldsymbol{\delta};\mathcal{F}\bigr{)}\triangleq J\!\left(\mathbf{x}+\boldsymbol{\delta},\,y;\mathcal{F}\right)\;-\;J\!\left(\mathbf{x},\,y;\mathcal{F}\right),$
where $J$ is the loss function, and $\mathbf{x}+\boldsymbol{\delta}$ corresponds to the adversarial example $\mathbf{x}^{\prime}$ in (II-A).
Building on ALG, we define the dual relative and absolute transferability gaps (rATG and aATG) to quantify cross-model transferability (Definition 3). The rATG quantifies transferability by contrasting cross-model ALG differences, where positive values indicate enhanced perturbation efficacy on the target model. Based on rATG, aATG incorporates the natural loss discrepancy $J(\mathbf{x},y;F^{\prime})-J(\mathbf{x},y;F)$, explicitly addressing inherent model divergence on clean inputs. This makes aATG more practical for real-world scenarios where models exhibit heterogeneous behaviors even without perturbations.
Definition 3 (Relative and Absolute Transferability Gap).
Given two models $\mathcal{F}$ and $\mathcal{F^{\prime}}$, an input–label pair $(\mathbf{x},y)$, and an adversarial perturbation $\boldsymbol{\delta}$, define
$\begin{split}\operatorname{rATG}\bigl{(}(\mathbf{x},y),\boldsymbol{\delta};\mathcal{F},\mathcal{F^{\prime}}\bigr{)}&\triangleq\operatorname{ALG}\bigl{(}(\mathbf{x},y),\boldsymbol{\delta};\mathcal{F^{\prime}}\bigr{)}\\ &\quad-\operatorname{ALG}\bigl{(}(\mathbf{x},y),\boldsymbol{\delta};\mathcal{F}\bigr{)},\end{split}$
the relative adversarial transferability gap. Then the absolute adversarial transferability gap is
$\begin{split}\operatorname{aATG}\bigl{(}(\mathbf{x},y),\boldsymbol{\delta};\mathcal{F},\mathcal{F^{\prime}}\bigr{)}\triangleq\;&\operatorname{rATG}\bigl{(}(\mathbf{x},y),\boldsymbol{\delta};\mathcal{F},\mathcal{F^{\prime}}\bigr{)}\\ &+\Bigl{[}J\!\left(\mathbf{x},\,y;\mathcal{F^{\prime}}\right)\;-\;J\!\left(\mathbf{x},\,y;\mathcal{F}\right)\Bigr{]}.\end{split}$
Notably, both rATG and aATG typically yield negative values in practical attacks: while adversarial perturbations $\boldsymbol{\delta}$ maximize $J(\mathbf{x}+\boldsymbol{\delta},y;\mathcal{F})$ on the surrogate model $\mathcal{F}$, they often induce smaller loss on $\mathcal{F^{\prime}}$, making positive transferability gaps unrealistic. Crucially, aATG simplifies to:
$\operatorname{aATG}\bigl{(}(\mathbf{x},y),\boldsymbol{\delta};\mathcal{F},\mathcal{F^{\prime}}\bigr{)}=J(\mathbf{x}+\boldsymbol{\delta},y;\mathcal{F^{\prime}})-J(\mathbf{x}+\boldsymbol{\delta},y;\mathcal{F}).$
When $\operatorname{aATG}<0$, successful transfer attacks require: minimizing $|\operatorname{aATG}|$ to align adversarial effects across models, and maximizing $J(\mathbf{x}+\boldsymbol{\delta},y;\mathcal{F})$ to elevate the upper bound of $J(\mathbf{x}+\boldsymbol{\delta},y;\mathcal{F^{\prime}})$. This joint strategy ensures sufficient adversarial impact on $\mathcal{F^{\prime}}$ despite negative gaps, formally linking to our subsequent flatness analysis.
In the next section, we will rigorously characterize the relationship between adversarial transferability and loss landscape flatness by leveraging the proposed absolute transferability gap (aATG) and average-case flatness ($\overline{R}^{(n)}_{\xi}$).
@@ IV-B Flatness and Transferability
Based on the analysis established in Section IV-A, achieving successful adversarial transfer requires simultaneously addressing two critical objectives: (1) maximizing the adversarial loss $J(\mathbf{x}+\boldsymbol{\delta},y;\mathcal{F})$ on the surrogate model $\mathcal{F}$ to amplify perturbation efficacy, and (2) minimizing the absolute adversarial transferability gap $|\operatorname{aATG}|$ to align the adversarial effects between the surrogate model $\mathcal{F}$ and the target model $\mathcal{F^{\prime}}$. While the first objective is straightforwardly addressed through gradient-ascent-based attack methods (e.g., PGD [53]), the second objective, systematically reducing $|\operatorname{aATG}|$, remains theoretically understudied and practically challenging. This constitutes the central focus of our theoretical analysis.
The minimization of $|\operatorname{aATG}|$ hinges on characterizing the geometric alignment of adversarial perturbations between divergent loss landscapes. We address this by establishing a multi-order flatness-dependent bound on $|\operatorname{aATG}|$ through Theorem 1, which directly connects transferability gaps to three measurable geometric properties: surrogate model flatness $\bar{R}^{(n)}_{\xi}(\mathbf{x};F)$, Target model flatness $\bar{R}^{(n)}_{\xi}(\mathbf{x};F^{\prime})$, and Cross-model gradient discrepancy $C_{n}(\mathbf{x})$.
@@
Let $\mathcal{F}$ and $\mathcal{F^{\prime}}$ be two models with associated loss $J(\mathbf{x},y;\,\cdot\,)$, and fix a point $(\mathbf{x},y)$. For each integer $n\geq 0$, define
$C_{n}(\mathbf{x})=\mathbb{E}_{\mathbf{x^{\prime}}\sim\mathrm{Unif}(B_{\xi}(\mathbf{x}))}\|\nabla^{n}J(\mathbf{x^{\prime}},y;F^{\prime})-\nabla^{n}J(\mathbf{x^{\prime}},y;F)\|,$
@@
and let $\bar{R}^{(n)}_{\xi}(\mathbf{x};F)$ and $\bar{R}^{(n)}_{\xi}(\mathbf{x};F^{\prime})$ be the $\xi$-radius $n$-order flatness for $F$ and $F^{\prime}$ respectively. Then for any perturbation $\boldsymbol{\delta}$ with $\|\boldsymbol{\delta}\|\leq\xi$, the absolute adversarial transferability gap satisfies
@@
$\begin{split}\bigl{|}\mathrm{aATG}((\mathbf{x},y),\boldsymbol{\delta};\mathcal{F},F^{\prime})\bigr{|}&\leq\sum_{n=0}^{\infty}\frac{\|\boldsymbol{\delta}\|^{n}}{n!}\Bigl{[}\bar{R}^{(n)}_{\xi}(\mathbf{x};F)\\ &\quad+C_{n}(\mathbf{x})+\bar{R}^{(n)}_{\xi}(\mathbf{x};F^{\prime})\Bigr{]}.\end{split}$
The proof of Theorem 1 (Proof IV-B) leverages multivariate Taylor expansions to decompose $\operatorname{aATG}$ into infinite-order curvature alignment terms, followed by norm-based bounding of each order’s contribution. This explicitly reveals how suppressing multi-order flatness variations ($n\geq 0$) and gradient discrepancies ($C_{n}$) jointly minimize transferability gaps.
Proof of Theorem 1.
First, by applying the multivariate Taylor expansion of the loss around $\mathbf{x}$ for both $\mathcal{F^{\prime}}$ and $\mathcal{F}$, we have
$J(\mathbf{x}+\boldsymbol{\delta},y;F^{\prime})$ $=J(\mathbf{x},y;F^{\prime})+\sum_{n=1}^{\infty}\frac{1}{n!}\,\nabla^{n}_{\mathbf{x}}J(\mathbf{x},y;F^{\prime})\bigl{[}\boldsymbol{\delta}^{\otimes n}\bigr{]},$
$J(\mathbf{x}+\boldsymbol{\delta},y;F)$ $=J(\mathbf{x},y;F)+\sum_{n=1}^{\infty}\frac{1}{n!}\,\nabla^{n}_{\mathbf{x}}J(\mathbf{x},y;F)\bigl{[}\boldsymbol{\delta}^{\otimes n}\bigr{]}.$
Subtracting these two series and including the zeroth‐order term yields
$\begin{split}\operatorname{aATG}\bigl{(}(\mathbf{x},y),\boldsymbol{\delta};\mathcal{F},\mathcal{F^{\prime}}\bigr{)}&=\sum_{n=0}^{\infty}\frac{1}{n!}\Bigl{(}\nabla^{n}_{\mathbf{x}}J(\mathbf{x},y;F^{\prime})\\ &-\nabla^{n}_{\mathbf{x}}J(\mathbf{x},y;F)\Bigr{)}\bigl{[}\boldsymbol{\delta}^{\otimes n}\bigr{]}.\end{split}$
For each $n\geq 0$, set
$T_{n}(\mathbf{x})=\nabla^{n}_{\mathbf{x}}J(\mathbf{x},y;F^{\prime})-\nabla^{n}_{\mathbf{x}}J(\mathbf{x},y;F).$
Then for any $\mathbf{x^{\prime}}\in B_{\xi}(\mathbf{x})$ one checks
$T_{n}(\mathbf{x})$ $=\bigl{[}\nabla^{n}J(\mathbf{x^{\prime}},y;F^{\prime})-\nabla^{n}J(\mathbf{x},y;F^{\prime})\bigr{]}$
$\quad+\bigl{[}\nabla^{n}J(\mathbf{x^{\prime}},y;F^{\prime})-\nabla^{n}J(\mathbf{x^{\prime}},y;F)\bigr{]}$
$\quad+\bigl{[}\nabla^{n}J(\mathbf{x^{\prime}},y;F)-\nabla^{n}J(\mathbf{x},y;F)\bigr{]}.$
Taking norms, applying the triangle inequality, and averaging over $\mathbf{x^{\prime}}\sim B_{\xi}(\mathbf{x})$ gives
$\|T_{n}(\mathbf{x})\|\;\leq\;\underbrace{\mathbb{E}_{\mathbf{x^{\prime}}\sim\mathrm{Unif}(B_{\xi}(\mathbf{x}))}\|\nabla^{n}J(\mathbf{x^{\prime}},y;F^{\prime})-\nabla^{n}J(\mathbf{x^{\prime}},y;F)\|}_{C_{n}(\mathbf{x})}\\ \;+\;\underbrace{\mathbb{E}_{\mathbf{x^{\prime}}\sim\mathrm{Unif}(B_{\xi}(\mathbf{x}))}\|\nabla^{n}J(\mathbf{x^{\prime}},y;F)-\nabla^{n}J(\mathbf{x},y;F)\|}_{\bar{R}^{(n)}_{\xi}(\mathbf{x};F)}\\ \;+\;\underbrace{\mathbb{E}_{\mathbf{x^{\prime}}\sim\mathrm{Unif}(B_{\xi}(\mathbf{x}))}\|\nabla^{n}J(\mathbf{x^{\prime}},y;F^{\prime})-\nabla^{n}J(\mathbf{x},y;F^{\prime})\|}_{\bar{R}^{(n)}_{\xi}(\mathbf{x};F^{\prime})}.$
Finally, since $\bigl{|}T_{n}(\mathbf{x})[\boldsymbol{\delta}^{\otimes n}]\bigr{|}\leq\|T_{n}(\mathbf{x})\|\|\boldsymbol{\delta}\|^{n}$, we get
$\begin{split}\bigl{|}\mathrm{aATG}((\mathbf{x},y),\boldsymbol{\delta};\mathcal{F},F^{\prime})\bigr{|}&\leq\sum_{n=0}^{\infty}\frac{\|\boldsymbol{\delta}\|^{n}}{n!}\Bigl{[}\bar{R}^{(n)}_{\xi}(\mathbf{x};F)\\ &\quad+C_{n}(\mathbf{x})+\bar{R}^{(n)}_{\xi}(\mathbf{x};F^{\prime})\Bigr{]},\end{split}$
as claimed. ∎
In transfer-based adversarial attacks where only the surrogate model $\mathcal{F}$ is accessible, the controllable parameter reduces to optimizing the source flatness $\bar{R}^{(n)}_{\xi}(\mathbf{x};F)$, as target model flatness $\bar{R}^{(n)}_{\xi}(\mathbf{x};F^{\prime})$ and cross-model discrepancy $C_{n}(\mathbf{x})$ remain intrinsically unmanageable. While TPA [2] attempted to theoretically link flatness and transferability, their analysis only considered first-order flatness ($n=1$) without incorporating higher-order gradient interactions ($n\geq 2$), leading to the erroneous conclusion that flatness minimally impacts transfer performance. Our theory thus formalizes why existing flatness regularization heuristics [7, 5, 8, 3, 6] succeed without target model knowledge, establishing flatness minimization on the surrogate model as both necessary and sufficient for practical transferability enhancement.
@@ V Method
Building on the theoretical linkage between multi-order flatness and transferability established in Section IV, we now address the second challenge outlined in Section III-C: designing a principled mechanism to balance local exploitation and global exploration. Motivated by Theorem 1’s identification of zeroth-order flatness dominance under bounded perturbations, our framework strategically focuses on optimizing $\overline{R}^{(0)}_{\xi}(\mathbf{x})$ through maximin expected flatness, neighborhood conditional sampling, and gradient balancing optimization to resolve the exploitation-exploration dilemma.
@@ V-A Maximin Expected Flatness
$\widehat{\overline{R}}^{(0)}_{\gamma,\xi}(\mathbf{x})\triangleq\max_{\mathbf{x}^{\rm c}\in B_{\gamma}(\mathbf{x})}\overline{R}^{(0)}_{\xi}(\mathbf{x}^{\rm c}),$
@@
With the proposed $\widehat{\overline{R}}^{(0)}_{\gamma,\xi}$, we establish the adversarial example generation framework guided by Theorem 1. As analyzed in Section IV-A, successful transfer attacks require: (i) minimizing $|\operatorname{aATG}|$ to align adversarial effects across models, and (ii) maximizing $J(\mathbf{x}+\boldsymbol{\delta},y;\mathcal{F})$ to elevate the upper bound of target model loss $J(\mathbf{x}+\boldsymbol{\delta},y;\mathcal{F^{\prime}})$. To unify these objectives, we formulate the attack generation as a multi-objective optimization problem:
@@
Substituting Definition V-A into (1) yields:
[…]
@@
Algorithm 1 Neighborhood Conditional Sampling
1:Surrogate model $\mathcal{F}$, loss function $J$; a raw example $x$ with ground-truth label $y$; neighborhood radius $\xi$, exploration radius $\gamma$; the number of sampling points $N$ and the max number of sampling iteration $T$
2:A set of sampling points $\{\mathbf{x}_{i}\}_{i=1}^{N}$
3:# Random initialization
4:Initialize ${\{\mathbf{x}_{0,i}\}}_{i=1}^{N}\in U_{\gamma}(\mathbf{x})$
5:for $t=0,1,\ldots,T-1$ do
6:  # Calculate the gradient of $J$
7:  $\{g_{t,i}\}_{i=1}^{N}=\nabla_{x}J({\{\mathbf{x}_{t,i}\}}_{i=1}^{N},y;\mathcal{F})$
8:  # Update the points in sub-region
9:  ${\{\mathbf{x}_{t+1,i}\}}_{i=1}^{N}={\{\mathbf{x}_{t,i}\}}_{i=1}^{N}-\xi\cdot sign(\{g_{t,i}\}_{i=1}^{N})$
10:  ${\{\mathbf{x}_{t+1,i}\}}_{i=1}^{N}=\prod_{{B}_{\gamma}(\mathbf{x})}(\prod_{{B}_{\xi}(\mathbf{x}_{0,i})}(\{\mathbf{x}_{t+1,i}\}_{i=1}^{N}))$
11:end for
[…]
13:return $\{\mathbf{x}_{i}\}_{i=1}^{N}$

### v3

@@ III-C Design Challenges
Existing flatness-enhanced methods face two unresolved challenges: theoretical ambiguity and optimization imbalance. First, the heuristic adoption of flatness metrics lacks rigorous theoretical grounding to connect geometric flatness with adversarial transferability [7, 8], with some work even questioning their inherent correlation [2], leading to inconsistent metric selection and suboptimal performance. Second, as shown in Section III-B, current methods suffer from a critical exploitation-exploration trade-off: average-case flatness optimization over-exploits local gradient consensus, causing premature convergence to suboptimal plateaus, while worst-case criteria induce over-exploration of non-informative sensitivity peaks, resulting in oscillatory updates. Addressing these challenges requires developing a unified framework that integrates theoretical insights into flatness-transferability relationships with adaptive mechanisms to harmonize local exploitation and global exploration, thereby maximizing adversarial transferability.
@@ IV Theoretical Analysis
Having formalized flatness metrics (Section III), we now resolve the first key challenge: the ambiguous relationship between geometric flatness and adversarial transferability. We first formally define transferability via the adversarial loss gap, then prove that improved multi-order flatness enhances it. This theory not only addresses the heuristic limitations of prior work but also provides principled optimization criteria.
@@ IV-A Adversarial Transferability Formulation
To formally characterize transferability, we propose the Adversarial Transferability Gap (ATG). We premise this definition on three explicit assumptions regarding the surrogate $\mathcal{F}$ and target $\mathcal{F^{\prime}}$: (1) they share the same label space; (2) they employ the same loss function $J(\cdot)$; and (3) they exhibit aligned performance on clean samples (i.e., $J(\mathbf{x},y;\mathcal{F})\approx J(\mathbf{x},y;\mathcal{F^{\prime}})$).
Definition 2 (Adversarial Transferability Gap).
Given the aligned models $\mathcal{F}$ and $\mathcal{F^{\prime}}$ (as assumed above), the ATG for an input–label pair $(\mathbf{x},y)$ and perturbation $\boldsymbol{\delta}$ is defined as:
$\operatorname{ATG}\bigl((\mathbf{x},y),\boldsymbol{\delta};\mathcal{F},\mathcal{F^{\prime}}\bigr)\triangleq J(\mathbf{x}+\boldsymbol{\delta},y;\mathcal{F^{\prime}})-J(\mathbf{x}+\boldsymbol{\delta},y;\mathcal{F}).$ (4)
Remark A detailed justification of this metric’s reasonableness and a discussion of its limitations under assumption violations are provided in Section -A of the Supplementary Material.
[…]
@@ IV-B Flatness and Transferability
Based on the definition established in Section IV-A, particularly in the typical scenario where transferability degrades (i.e., $ATG<0$), achieving successful adversarial transfer requires simultaneously addressing two critical objectives: (1) maximizing the adversarial loss $J(\mathbf{x}+\boldsymbol{\delta},y;\mathcal{F})$ on the surrogate model $\mathcal{F}$ to strengthen the attack,, and (2) minimizing $|\operatorname{ATG}|$ to align adversarial effects between the surrogate $\mathcal{F}$ and the target $\mathcal{F^{\prime}}$. While gradient-ascent attacks (e.g., PGD [43]) readily achieve the first, systematically reducing $|\operatorname{ATG}|$ remains theoretically understudied and challenging. We therefore make this our theoretical focus.
The minimization of $|\operatorname{ATG}|$ hinges on characterizing the geometric alignment of adversarial perturbations between divergent loss landscapes. We address this by establishing a multi-order flatness-dependent bound on $|\operatorname{ATG}|$ through Theorem 1, which directly connects transferability gaps to three measurable geometric properties: surrogate model flatness $\bar{R}^{(n)}_{\xi}(\mathbf{x};F)$, Target model flatness $\bar{R}^{(n)}_{\xi}(\mathbf{x};F^{\prime})$, and Cross-model gradient discrepancy $C_{n}(\mathbf{x})$.
@@
Let $\mathcal{F}$ and $\mathcal{F^{\prime}}$ be two models with loss $J(\mathbf{x},y;\,\cdot\,)$. Assume $J$ is locally approximable by a Taylor expansion of order $N$. For $0\leq n\leq N$, define
$C_{n}(\mathbf{x})=\mathbb{E}_{\mathbf{x^{\prime}}\sim\mathrm{Unif}(B_{\xi}(\mathbf{x}))}\|\nabla^{n}J(\mathbf{x^{\prime}},y;F^{\prime})-\nabla^{n}J(\mathbf{x^{\prime}},y;F)\|,$ (5)
@@
and let $\bar{R}^{(n)}_{\xi}(\mathbf{x};F)$ and $\bar{R}^{(n)}_{\xi}(\mathbf{x};F^{\prime})$ be the $n$-order flatness for $F$ and $F^{\prime}$. Then for perturbation $\|\boldsymbol{\delta}\|\leq\xi$, the adversarial transferability gap satisfies
@@
$\begin{split}\bigl|\mathrm{ATG}((\mathbf{x},y),\boldsymbol{\delta};\mathcal{F},F^{\prime})\bigr|&\leq\sum_{n=0}^{N}\frac{\|\boldsymbol{\delta}\|^{n}}{n!}\Bigl[\bar{R}^{(n)}_{\xi}(\mathbf{x};F)\\ &\qquad+C_{n}(\mathbf{x})+\bar{R}^{(n)}_{\xi}(\mathbf{x};F^{\prime})\Bigr]\\ &\qquad+\Delta R_{N}(\mathbf{x},\boldsymbol{\delta}),\end{split}$ (6)
where $\Delta R_{N}$ bounds the difference in higher-order residuals.
The proof of Theorem 1 (see Section -B in Supplementary Material) employs finite-order multivariate Taylor expansions to decompose $\operatorname{ATG}$ into curvature alignment terms plus a residual bound. This framework reveals that suppressing multi-order flatness variations and gradient discrepancies ($C_{n}$) minimizes the transferability gap. Crucially, $C_{n}(\mathbf{x})$ quantifies intrinsic surrogate-target dissimilarity; a larger $C_{n}$ loosens the bound, acting as a natural barrier to transfer attack success.
In transfer-based attacks where only surrogate $\mathcal{F}$ is accessible, optimizing source flatness $\bar{R}^{(n)}_{\xi}(\mathbf{x};F)$ becomes the sole effective strategy, as target flatness and $C_{n}(\mathbf{x})$ are uncontrollable. While TPA [2] linked flatness to transferability, it was limited to first-order analysis ($n=1$). Our theory incorporates higher-order interactions ($n\geq 1$), formalizing why heuristics [7, 5, 8, 3, 6] succeed. Theorem 1 also shows that high-order terms diminish rapidly due to the $\frac{\|\boldsymbol{\delta}\|^{n}}{n!}$ coefficient. For ReLU networks where high-order derivatives vanish ($N\approx 1$), the zeroth-order flatness $\bar{R}^{(0)}_{\xi}(\mathbf{x};F)$ becomes the dominant factor, establishing it as an efficient optimization target.
@@ V Method
(a) The Workflow of MEF Optimization
(b) Geometric Interpretation
Figure 3: Overview of the proposed MEF attack. (a) Illustration of the iterative optimization process. At step t1t-1, Neighborhood Conditional Sampling (NCS) transforms random samples ($x^{r}$) into worst-case neighborhood samples ($x^{n}$) via inner minimization. Outer maximization then aggregates gradients from these regions to update the adversarial example $x_{t}^{adv}$. (b) Geometric interpretation showing how MEF seeks flat maxima to minimize the Adversarial Transferability Gap (ATG) between the source model $F$ (solid line) and the target model $F^{\prime}$ (dashed line), as opposed to sharp maxima which lead to larger gaps.
Building on the theoretical linkage between multi-order flatness and transferability established in Section IV, we now address the second challenge outlined in Section III-C: designing a principled mechanism to balance local exploitation and global exploration. Instead of heuristically adopting existing metrics, our framework is strictly guided by Theorem 1: since the impact of higher-order terms ($n\geq 1$) diminishes exponentially due to the $\frac{\|\delta\|^{n}}{n!}$ coefficient, we prioritize the theoretically dominant zeroth-order flatness. Crucially, to resolve the exploitation-exploration dilemma inherent in standard definitions, we define a novel Zeroth-order Worst-Neighborhood Average Flatness and optimize it through maximin expected flatness, neighborhood conditional sampling, and gradient balancing optimization.
@@ V-A Maximin Expected Flatness
$\widehat{\overline{R}}^{(0)}_{\gamma,\xi}(\mathbf{x})\triangleq\max_{\mathbf{x}^{\rm c}\in B_{\gamma}(\mathbf{x})}\overline{R}^{(0)}_{\xi}(\mathbf{x}^{\rm c}),$ (7)
@@
With the proposed $\widehat{\overline{R}}^{(0)}_{\gamma,\xi}$, we establish the adversarial example generation framework guided by Theorem 1. As analyzed in Section IV-A, successful transfer attacks require: (i) minimizing $|\operatorname{ATG}|$ to align adversarial effects across models, and (ii) maximizing $J(\mathbf{x}+\boldsymbol{\delta},y;\mathcal{F})$ to elevate the upper bound of target model loss $J(\mathbf{x}+\boldsymbol{\delta},y;\mathcal{F^{\prime}})$. To unify these objectives, we formulate the attack generation as a multi-objective optimization problem:
@@
Substituting Definition 7 into (8) yields:
$\min_{\mathbf{x}^{\rm adv}\in B_{\epsilon}(\mathbf{x})}\Bigl\{$ $\max_{\mathbf{x}^{\rm c}\in B_{\gamma}(\mathbf{x}^{\rm adv})}\mathbb{E}_{\mathbf{x}^{\prime}\sim\mathrm{Unif}(B_{\xi}(\mathbf{x}^{\rm c}))}\bigl[\bigl|J(\mathbf{x}^{\prime})-J(\mathbf{x}^{\rm c})\bigr|\bigr]$